OpenStudy (anonymous):

lambda^3 - 3lambda^2-18 + 0 = 0 If possible could someone show me how to find the roots of this equation, step by step, thanks.

OpenStudy (anonymous):

does that really say +0= 0 at the end?

OpenStudy (anonymous):

you can drop it off i guess

OpenStudy (anonymous):

so you are solving \[x^3-3x^2-18=0\] use a calculator because you are not going to do this by hand http://www.wolframalpha.com/input/?i=x%5E3-3x%5E2-18%3D0

OpenStudy (anonymous):

i understand the roots are lambda = -3 and lambda = 6 just not sure how to arrive at that answer with all working

OpenStudy (anonymous):

need to show working it's for an assignment

OpenStudy (anonymous):

then as you can see by the link i sent the equation is wrong or the solutions are wrong

OpenStudy (anonymous):

now if the equation was \[\lambda^2-3\lambda-18=0\] then the solutions would in fact be \[\lambda = -3,\lambda =6\]

OpenStudy (anonymous):

yes thats correct, ow do you get -3 and 6

OpenStudy (anonymous):

factor \[(\lambda +3)(\lambda -6)=0\] etc

OpenStudy (anonymous):

oooooooooooooooh you have a different equation!

OpenStudy (anonymous):

you have \[\lambda ^3 -3\lambda ^2-18\lambda =0\]

OpenStudy (anonymous):

yep, i take a lambda outside to get your equation though cant i?

OpenStudy (anonymous):

make a big difference. in that case you still factor as \[\lambda(\lambda+3)(\lambda -6)=0\] and get three solutions. -3,0,6 your question was missing the last lambda

OpenStudy (anonymous):

yes, i have it now thanks satelite73